OpenStudy (anonymous):

Simplify each expression. Rationalize all denominators. Assume that all variables are positive. 1) square root of 200x^3y over square root of 2xy^5 2) (8-3square root of 2)(8+3square root of 2) Please explain and show steps!

OpenStudy (campbell_st):

Ist the numerator \[\sqrt{200x^3y} = \sqrt{2 \times 100 \times x^2 \times x \times y} = \sqrt{100x^2}\times \sqrt{2xy}\] this simplifes to \[10x \sqrt{2xy}\] now the denominator \[\sqrt{2xy^5} = \sqrt{y^4} \times \sqrt{2xy}\] which simplifes to \[y^2 \sqrt{2xy}\] the problem is now \[(10x \sqrt{2xy})/(y^2 \sqrt{2xy})\] cancelling common factors gives \[10x/y^2\]

OpenStudy (anonymous):

what does Rationalize all denominators mean?

OpenStudy (anonymous):

@campbell_st do you know how to do (8-3square root of 2)(8+3square root of 2)

OpenStudy (campbell_st):

This ia a binomial expansion called the difference of two squares \[(8 - 3\sqrt{2})(8 + 3\sqrt{2})\] then answer is \[8^2 - (3\sqrt{2})^2 = 64 - 9\times2\] I'll let you finish it